# TsCA: On the Semantic Consistency Alignment via Conditional Transport for Compositional Zero-Shot Learning

### v1

@@ Introduction
We compare our TsCA with the most recent CZSL methods. The results are shown in Tab. 2 and Tab 3. On the closed-world setting, TsCA exceeds the previous SOTA methods on all datasets in core metrics. Specifically, it yields improvements of 0.6% on MIT-States, 1.9% on UT-Zappos, and 0.5% on C-GQA in H over the second-best methods. It also attains considerable gains in AUC with increases of 0.5%, 0.7%, and 0.2%, respectively. Notably, the advantage of UT-Zappos is more significant, suggesting that the fine-grained nature of UT-Zappos can better align with TsCA’s ability to achieve precise alignment across local visual features and textual representations. Similarly, the results in the open-world setting are also promising. Our TsCA attains the highest AUC across all datasets, with the scores of 8.7% ,36.7% and 4.5%. It should be noted that the CT-based filtering strategy contributes to the improvements in this challenging setting. All numerical results substantiate our motivation to empower the model to capture semantic consistency among the image-composition-primitive interactions.
Table 4: Ablation results for each component on UT-Zappos in the closed-world setting.
$\mathcal{L}_{de}$ $\text{CT}_{{\bf C}}$ $\mathcal{L}_{cyc}$ S U H AUC
1 65.7 73.5 54.4 40.8
2 68.2 73.9 56.8 44.3
3 67.0 74.1 56.7 43.8
4 67.7 74.4 57.3 44.6
5 68.5 74.5 57.6 44.9
6 68.6 74.3 58.4 45.1
@@
Conclusion
In this paper, we revisit compositional generalization in CZSL through a conditional transport perspective. We explore pairwise CTs between the local visual features of images and two textual label sets within a multi-path paradigm. We then introduce cycle-consistency as a link to bond all sets, promoting robust learning. The primitive decoupler further improved accuracy and decoupled global primitive visual representations during prediction. Benefiting from the transport plan between the primitive set and the composition set, this approach successfully narrowed the composition search space by excluding unfeasible pairs during inference. Extensive experiments across three benchmarks consistently demonstrate the superiority of TsCA. Comprehensive ablation studies and visualizations confirm our motivation and the essential role of each component. With its inherent flexibility and simplicity, we hope our work inspires innovative ideas for future research.

### v3

@@ 4.2 Comparision with State-of-the-Arts
We compare our TsCA with the most recent CZSL methods. The results are shown in Tab. 1 and Tab. 2. On the closed-world setting, TsCA exceeds the previous SOTA methods on all datasets in core metrics. Specifically, it yields improvements of 0.6% on MIT-States, 2% on UT-Zappos, and 1.1% on C-GQA in H over the second-best methods. It also attains considerable gains in AUC with increases of 0.5%, 1.6%, and 0.8%, respectively. Notably, the advantage of UT-Zappos is more significant, suggesting that the fine-grained nature of UT-Zappos can better align with TsCA’s ability to achieve precise alignment across local visual features and textual representations. Similarly, the results in the open-world setting are also promising. Our TsCA attains the highest AUC across all datasets , with the scores of 8.7% , 37.1% and 4.5%. It should be noted that the CT-based filtering strategy contributes to the improvements in this challenging setting. All numerical results substantiate our motivation to empower the model to capture semantic consistency among the image-composition-primitive interactions.
Components closed-world open-world
S U H AUC S U H AUC
$\mathcal{L}_{base}$ 65.7 73.5 54.4 40.8 64.3 62.5 48.0 33.3
+ $\mathcal{L}_{de}$ 68.2 73.9 56.8 44.3 66.9 62.2 49.2 34.6
+ $\mathcal{CT}$ 67.8 74.3 57.5 44.6 67.0 64.6 50.1 35.9
+ $\mathcal{CT}$ + $\mathcal{L}_{cyc}$ 68.9 74.5 58.0 45.0 67.6 63.6 51.9 36.2
+ $\mathcal{L}_{de}$ + $\mathcal{CT}$ 69.0 75.4 58.2 45.3 67.9 63.7 52.0 36.3
+ $\mathcal{L}_{de}$ + $\mathcal{CT}$ + $\mathcal{L}_{cyc}$ 68.7 75.8 58.5 46.1 69.8 63.4 52.2 37.1
Table 3: Ablation results for each component on UT-Zappos.
@@ 4.4 Qualitative Results
5 Conclusion
In this paper, we revisit compositional generalization in CZSL through a conditional transport perspective. We explore pairwise CTs between the local visual features of images and two textual label sets within a multi-path paradigm. We then introduce cycle-consistency as a link to bond all sets, promoting robust learning. The primitive decoupler further improved accuracy and decoupled global primitive visual representations during prediction. Benefited from transport plan between primitive set and composition set, this approach successfully narrowed the composition search space by excluding unfeasible pairs during inference. Extensive experiments across three benchmarks consistently demonstrate the superiority of TsCA. Comprehensive ablation studies and visualizations confirm our motivation and the essential role of each component. With its inherent flexibility and simplicity, we hope our work inspires innovative ideas for future research.